AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 12 of 33  ELOG logo
  ID Date Author Subject
  445   Tue Sep 26 10:13:14 2023 TDELOTECH PID Temperature Controller R4000 manual ( German and English )
  444   Tue Sep 26 08:08:42 2023 TDTuesday 26 September
09.07 IE514 pressure 6.4e-11 mbar, ambient temperature +20.3 deg C

      Using appc218 ( account carme ) workstation by the CARME 19" rack 
     
      ssh pi@carmebakingpi

      Re-connect to bakeout monitor app 

      screen -dR

      See attcahment 1


12.02 Vacuum group ion pump disconnected and RGA removed from CARME chamber -  outside of CARME bakeout tent 'closed' - see attachments 2-4

      Vacuum group IE514 to be disconnected


14.45 Vacuum group disconnect their IE514

      Inside of CARME bakeout tent closed

      Photos of the CARME bakeout tent inside and outside CRYRING - see attachments 5-7  
  443   Mon Sep 25 16:33:32 2023 TDRemoved tape from heaters TC
17.33 IE514 pressure 6.4e-11 mbar, ambient temperature +20.8 deg C

      The thermocouples on the 3x groups of heater elements have been detached from direct contact with the heating element and now float abpove the heating element by 2-
3cm - see attachments 1 & 2

      The vacuum group IE514 and Ametek RGA are still attached to CARME - see attachment 3. Have requested vacuum group remove them.

      The 'air side' of the Be window on the interaction chamber is under vacuum provided by small Agilent scroll pump.

      
  442   Fri Sep 22 12:35:46 2023 JMBakeout plots

Attached are plots for the previous bakeout of CARME which may aid in the bakeout to come in the following week.  

Plots relate the heater temperatures and heater powers with the chamber temperature (internal floating thermocouple) and pressure inside CARME. 

  441   Wed Sep 13 15:00:39 2023 MLVacuum and Heater temperature plots from Grafana

Attached are the plotted curves from vacuum and heater temperature (actual and set values from ISOHEAT controllers) during the bakeout process. Channel assignment as per post 433.

  440   Sat Sep 2 10:04:55 2023 CBTurbo GV closed

Found CARME in XHV at 1.1E-11 mbar. Chamber temperature around 30C cooling slowly.

Closed GV to top turbo. Pressure goes up but appears to recover around 2.5E-11 mbar, on line with several previous observations that sectioning off the turbo results in a higher reading in the IE514.

Left turbo on. Remember to not turn it off when removing the tent.

  439   Fri Sep 1 08:11:40 2023 CBNEG activation - 2

9:10 Set heaters temperature 300 -> 250 C in preparation for cooling down and to help keep DSSD temperature down. Started activating set 6 (3 X DSSD) at 120 W as per previous procedure.

9:15 Realised previous choice is a mistake in the current configuration as the DSSD TC is closer to set 5. Set 6 off. Set 5 to 60W. No effects in temperature. Ramped slowly to 200W. No immediate effect in temperature, while vacuum goes up.

9:30 Started activating set 6 at 300W (=100 per module). This should give us the best feedback on DSSD temperature from the only TC we have since they should all be at similar temperatures. 

9:40 DSSD TC shows temperature decreasing rapidly (-17C/h). That's nonsense. The TC cannot be working properly, which means activation is now quite dangerous. Looking at strain relief TC instead from now, which is still at 85C. Set 5 down to 30W for safety while set 6 finishes ramping up.

9:45 DSSD TC still reading nonsense at 35C cooling by 20C per hour. Strain relief TC at 85C heating up by 10C per hour is more realistic. Vacuum worsening as set 6 reached 300W. Immediately ramped down to 120W. Leaving for 30 minutes, or less if strain relief temperature increases too much.

9:57 Strain relief heating quickly at 90C. Decided to turn off ALL negs for safety. Set heaters to 200C.

10:18 Strain relief temperature finally slowing down heating rate. DSSD TC beings displaying increasing temperature. The TC wire on the TC-08 module is clearly damaged.

 

After lunch, turned on the four ion elements of the D-2000. Pump#4 is not working, and apparently module #2 also has issues. Used module #4 on pump#2 to get three working pumps. Others will have to be investigated when the tent is removed and we can have some access to the pumps.

Tried closing gate valve to turbo. Vacuum does not seem to really improve. Opened. Will try again tomorrow morning.

Vacuum division ion pump on as well.

  438   Thu Aug 31 08:36:34 2023 CBNEG activation

Found CARME at 50C, a bit cooler than expected, probably due to invasive tent changes yesterday to mount missing wires. Increased heaters temperature to 300C

 

9:35 Started D-2000 #1,3 (behind the DSSD). T=400, tramp 30, thold 90

11:20 Set D-2000 1,3 to 200C. Turned on D-2000 2,4 as above 

13:15 Set D-2000 2,4 to 200C. Turned on UHV1400 set 1 at 250W, tramp 30, thold infinity (will ramp down manually as required)

13:20 No appreciable effects on DSSD temperature. Turned on set 2 as well, which should be mounted close nearby.

13:45 Both sets reached 250W. No effects on detector temperature. Increased to 300W. Vacuum gets slightly worse, but recovers quickly. Not worth further increasing power. Leaving for 90 minutes.

15:20 Strain relief TC at 75C quite hotter than detector TC at 35C. Might be normal. Vacuum division IE514 must have been degassed remotely? Decreased set 1,2 to 75W. Turned on sets 3,4 to 250W as above. 

15:45 Sets 3, 4 now at 250W. Detector TC heating quickly at 13C/h but nowhere near dangerous temperatures yet. Decreased sets 1,2 to 50W. Leaving for 90 minutes.

17:15 Sets 3, 4 to 50W. Slightly worrying temperature difference between detectors and strain relief around 65C. Different materials and different positions could explain this.

17:40 Oddly, detector temperature now going up sharply (temperature diff. down to 40C) since the uhv1400 power was decreased. Cannot think of s logical reason,  it may just be a coincidence and heat transfer is catching up.

 

Leaving overnight. Will turn on the last two groups of UHV1400 tomorrow morning.

  437   Wed Aug 30 10:28:36 2023 CBPreparing for pumps on

We need to connect the non-bakeable (max 80C) wire of the ion pump belonging to the vacuum division, and we also need to connect the two newly-repaired NEG cables to the D2000 modules. That requires partly opening the tent to gain access.

Set heaters temperature from 450 C -> 400 C - CARME floating TC temperature goes from ~98 C to ~88C.

Set heaters temperature to 350 C. Left to cool over lunch.

 

 Connected GSI vacuum division ion pump and IE514

Connected D-2000 upstream left/right with repaired cables

 

Flashed all ion pumps. Ion pump #4 was flashed a couple of times, but has now started arcing. The issue is not with the module - either the wire or the ion element are damaged.

Set heaters to 300 C. Left to cool a bit more overnight. Planning on NEG on tomorrow.

  436   Wed Aug 30 09:33:13 2023 CB, OGNEG cables repaired

Repaired Neg wires for D2000 with Oemer's help.

Cable #2 is fine. 

In Cable #4, wires 1 and 2 are swapped but it's impossible this happened when the connector came apart. The pins are crimped and secured with a screw. Most likely it was sold like this and there's no difference between wire 1 and 2 (same voltage).

Edit: SAES confirmed 1&2 are interchangeable.

  435   Mon Aug 14 09:22:56 2023 JMBakeout

14/08/23

Temperature in the baking volume is ~ 42 C and the pressure is 4.6e-8 mbar. Current heaters set point is 135 C. This will now be increased to 200 C to start increasing the baking temperature more significantly.

Some issues with the interaction chamber heaters were noticed this morning. This included very large increases in temperature despite only small increases in the setpoint. When this occurred the heater channels were simply turned off. Channel 3 on the interaction chamber (main jacket on the centre of the chamber was also observed to smoke slightly). This smoking was seen when the temperature for channel 3 was ~ 70 C. 

Changing the settings of these heaters appears to have resolved most of the issues. The heaters now ramp in a much more controlled manner and remain roughly constant at the setpoint value.  The interaction chamber heaters were all increased to 70 C with no issues observed. Channel 3 also appears OK.

The CARME heaters settings have also been changed and now the heaters temperature and power output do not oscillate as observed previously. Power output per heater is approximately 15 % of total at heater temperatures of 220 C, and this value remains consistent over time. The temperatures inside the baking volume are now ~ 50 C with the pressure equal to 7e-8 mbar. Heating has not been the priority while investigating the issues observed with the heaters.

As a precaution CARME will now be maintained at the 50C level. The interaction chamber and e-target temperatures have been set to 50 C also. If the system remains stable overnight we will again begin increasing the temperature in the morning.

15/08/23

Temperatures across the whole of YR09 have remained very stable overnight. Temperature of CARME has increased slightly and is now ~ 55 C. Pressure is ~1e-7 mbar. A brief lincrease in pressure was observed yesterday due to the de-gassing of the GSI ion pump on CARME. Plan is to now increase the temperatures of YR09 to the 100 C level 

Temperature is ~ 70 C. Heating is slower than expected as I am aiming to minimise thermal gradients across the chamber, which was previously done automatically by our baking code. Current heater temperature values are ~ 320 C. I have increased the temperatures across YR09 to 70 C to match carme . Pressure in CARME is 2.1e-7 mbar.

16/08/23 

The temperatures of the CARME heaters were further increased to ~ 420 C to further increase the chamber temperature. Average chamber temperatures of ~93C achieved. Heater temperatures can be pushed to 450C which should allow us to achieve 100 C. Will allow chamber to equillibriate as temperature gradient between top and bottom is quite large.

17/08/23  

Baking will be unsupervised over the weekend. To remain safe, the berylium window and the interaction chamber will be kept at a temperature of 50C. Carme will be reduced to ~60C during this period.

Attached are plots for the bakeout period concerning this week.  

  434   Sat Aug 12 15:06:51 2023 JMInitial bakeout tests

12/08/23

Baking testing. Aim to find relationship between heater temperature setpoint and the temperature of the chamber and floating temperature in the baking volume. Behaviour of the heater tempertures and ability to control the heating is also required. Ambient chamber temperature ~22 C. Interaction chamber temperatures will be raised to the level of the CARME chamber throughout, starting temperature of the interaction chamber with heating off is 28 C. P = 7.5e-9 mbar

CARME thermocouple temperatures can be read from the monitoring code which is currently running in a screen session on the baking pi. Heater temperatures and pressure in YR09 can be read from Grafana.

15:40 CARME heaters turned on, set point 50 C. Ramping is ~ 1C per minute for the heaters. 

16:00 Thermocouples inside baking volume show small increases. Setpoint increased to 60 C.

16:10 Setpoint increased to 80 C.

16:30 Setpoint increased to 100 C.

16:45 Setpoint increased to 120 C

17:00 Setpoint raised to 150 C. Average chamber temp ~ 25.2 C, bottom of the chamber is ~28C. 

17:45 Setpoint was maintained at 150 C, temperatures of the heaters continued to rise past this setpoint and were floating about 200-230 C with large variations in the temperature. At lower setpoints the heater temperature appeared to rise linearly with the 1 C per minute. I do not believe I currently have control of the heating. Average chamber temp 28.8 C with bottom of the chamber ~35-37 C. Heating has been turned off while I grab some food. P = 1e-8 mbar

Next test will be to try and maintain a consistent heater tempreture for a longer period of time, probabably at a lower set temperture. 

13/08/23

10:00 Heater setpoint to 50 C. Will maintain at this level for a while to look at increase in chamber temps and see if heater temp remains stable. P = 8.52e-9 mbar

11:47 Heater temperatures have been constant at ~ 55 C for the last hour. Little heating observed in the chamber, requires much larger temperature to heat the baking volume. Setpoint increased to 100 C.

13:10 Heater temperatures have been constant at ~103 C for the last  40 mins or so and appears stable. Temperatures are rising at a steady rate but temps are still low. Interaction chamber heating jackets activated to be a same temperature as CARME. Will require higher temperatures to heat the baking volume.

16:15 Heater temperaturtes have remained constant at ~103 C. Temps inside the baking volume remain low, Increasing heater temperature to 135 C.

18:30 Heaters temperatures remain stable. Larger fluctations are apparent than at a setpoint of 100C with temps between 140 and 150 C. Chamber temperatures increasing avg is ~ 35C, was aiming for 50C over the weekend. Increasing setpoints to 170 C.

20:50 Heater tempretures are stable. Average chamber temp is ~ 40 C. Interaction chamber heaters set to 40 C to keep constant with CARME. Pressure has increased to 2.9e-8 mbar. Heater power outputs are visible in last image, heaters currently turn on at ~80% max power about twice every 5 mins. I expect temps of 250-300 C are required to obtain significant temp increase up to 100 C on the chamber. Chamber temps appear to be fairly stable at this point, will leave for now and check back in a few hours.

23:00 Heater setpoint temp reduced to 135 C to maintain temperatures inside the baking volume overnight with no increase. Heater power has now reduced to < 50 % full power when the eaters are on. Pressure in the chamber is 4.4e-8 mbar.

I believe I now understand the heating system and plan to increase the chamber temperature throughout the day tomorrow. Aim is to get close to full bakeout temperatures.

  433   Sat Aug 12 11:57:51 2023 JMUpdated thermocouple positions and heater channel layouts

Heater channel layouts for the interaction chamber and carme in attached documents. 

Updated thermocouple positions; (front section is closest to gas-target, rear section is moving detector end)

  • 1-1 Internal TC 1
  • 1-2 Internal TC 2
  • 1-3 Internal TC 3
  • 1-4 CARME - top of the front section
  • 1-5 CARME - bottom of the front section
  • 1-6 CARME - top of the mid section
  • 1-7 CARME - middle of the mid section
  • 1-8 CARME - bottom of the mid section
  • 2-1 Float - bottom of mid section 
  • 2-2 CARME - middle of rear section 
  • 2-3 Float - middle of the front section
  • 2-4 CARME - bottom VAT valve
  • 2-5 Float - middle of the rear section 
  • 2-6 CARME - Top elbow
  • 2-7 Float - top of the mid section
  • 2-8 CARME - rear beam pipe

Thermocouple posiitons have been updated in the monitoring code, including the averages for the CARME chamber. 

  431   Thu Aug 10 07:51:43 2023 JMBaking tent and heater testing

09/08/23

Attached is the current pumpdown plot taken from grafana. 

I have started to re-work the baking tent in order to close gaps from its initial installation. The heaters work and were tested to a temperature of 50C. Testing is required to gauge the heater temperature required to heat the baking volume, we suspect the heaters will be several hundred degrees. 

11/08/23

Baking tent has been fully constructed again, see photos attached.  Some thermocouples have been re-positioned, new thermocouples positions are given in elog 433. GSI ion pump cable has been connected. Roughing pump on the outside of the berylium window has been connected and berylium window is at < 10 mbar. 

  430   Thu Jul 27 09:43:56 2023 RSSThermocouples connected -2
10:30  IE514 1.80e-08mbar

Most of the CARME is covered by the baking tent. The tent needs to be closed and a bit covered from the front (upstream), end (downstream), and bottom side. TCs are labeled with a marker in the sequence 1-1 to 2-8.

1-1 Internal TC 1
1-2 Internal TC 2
1-3 Internal TC 3
1-4 CARME front top
1-5 CARME front bottom
1-6 Top hanging middle
1-7 CARME middle middle
1-8 CARME middle bottom

2-1 CARME middle top
2-2 CARME back middle
2-3 Elbow
2-4 Valve
2-5 Magnet pipe
2-6 CARME back top
2-7 CARME back float
2-8 CARME back bottom

Also, TC extensions would be needed to connect to TC-8 modules.

  429   Wed Jul 26 08:28:09 2023 RSSBuilding tent
09:26  IE514 1.96e-08mbar
  428   Tue Jul 25 09:12:15 2023 RSSBuilding tent
10:10  IE514 2.0e-08mbar

Upstream part of the CARME is covered.

  427   Mon Jul 24 10:40:14 2023 RSS, TDPump down continues
11.40 Pirani 1.00e-3 mbar IE514 2.23e-08mbar

15.15 7d/24h vacuum plots
  426   Sun Jul 23 10:13:48 2023 RSS, TDBuilding tent
11.15 Pirani 1.0e-03mbar IE514 2.72e-08mbar

16.35 Pirani 1.0e-03mbar IE514 2.61e-08mbar
  425   Sat Jul 22 09:19:59 2023 RSS, TDThermocouples connected
10.20 Pirani 1e-03mbar IE514 3.26e-08mbar


12.53 CARME Leybold IE514 gauge OFF


Top right DSSSD bias -20V leakage current -10.51uA


Three internal ThermoCouples (TCs) -labeled 1, 2, and 3 in attachment 1 - were attached and tested on the TC-08 boards. Only 2 showed the ambient temperature, see attachment 3. 1 and 3 showed some strange numbers, see attachments 2 and 4 respectively.


15.28 CARME Leybold IE514 gauge OFF. Top right DSSSD bias -20V leakage current -10.44A


All the TCs are attached to TC-08 modules. See attachments 5-7. Once all TCs were plugged into TC-08 modules, they were fine.


TC-08 A0067

1. Internal TC 1
2. Internal TC 2
3. Internal TC 3
4. CARME front top
5. CARME front bottom
6. Top hanging middle
7. CARME middle middle
8. CARME middle bottom



TC-08 A0066

1. CARME middle top
2. CARME back middle
3. Elbow
4. Valve
5. Magnet pipe
6. CARME back top
7. CARME back float
8. CARME back bottom

Heating Jackets placed beneath the heaters. See attachments 8-10.
ELOG V3.1.4-unknown